OpenStudy (anonymous):

If the slope of the line that passes through the points (a,0) and (1,2) is 1/2, what is the value of a?

OpenStudy (anonymous):

draw a picture of a line through 1,2 with slope 1/2...

OpenStudy (anonymous):

...

OpenStudy (anonymous):

(y-2) = (1/2)*(x-1) also works

OpenStudy (anonymous):

(let y = 0 and solve that equation for x)

OpenStudy (anonymous):

Just use the formula to figure out slope: rise over run It's change in y over change in x \[(y2-y1)/(x2-x1)=(2-0)/(1-a)=2/(1-a)\]

OpenStudy (anonymous):

\[2/(1-a)=1/2\] \[1-a=1/4\] \[-a=-3/4\] a=3/4